OpenStudy (anonymous):

Could someone help me take the second derivative using implicit differentiation?

OpenStudy (anonymous):

I have a question and i did it, but im stuck on a part

OpenStudy (anonymous):

the first deri is -x^2/y^2

OpenStudy (phi):

Just to be sure we are on the right track, what is the original equation and question?

OpenStudy (anonymous):

x^3 +y^3 =16....find second deriv at (2,2)

OpenStudy (phi):

taking the derivative with respect to x. I do \[ \frac{d}{dx} \left( x^3 +y^3 =16\right)\\ \frac{d}{dx}x^3 + \frac{d}{dx}y^3 = \frac{d}{dx}16\\ 3x^2 \frac{dx}{dx}+3y^2 \frac{dy}{dx}=0\\ 3y^2 \frac{dy}{dx}= -3x^2\\ \frac{dy}{dx}= x^2 y^-2 \]

OpenStudy (phi):

so you did the first derivative correctly. now on to the second. \[ \frac{d^2y}{dx^2} = \frac{d}{dx}(x^2y^{-2} ) \\ = x^2 \frac{d}{dx}y^{-2} + y{-2}\frac{d}{dx}x^2 \] can you do the next steps

OpenStudy (anonymous):

can i show u what i did

OpenStudy (phi):

\[ \frac{d^2y}{dx^2} = \frac{d}{dx}(x^2y^{-2} ) \\ = x^2 \frac{d}{dx}y^{-2} + y^{-2}\frac{d}{dx}x^2 \]

OpenStudy (anonymous):

\[= \left(\begin{matrix}[-2x](y^2) - [2y dy/dex](-x^2) \\ y^4\end{matrix}\right)\]

OpenStudy (anonymous):

thats what i have i mess up when i sub in the first derive

OpenStudy (phi):

are you using the quotient rule ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

I think the signs are wrong. Try writing it as x^2 y^(-2) and use the product rule. we can come back to the quotient rule after we get the right answer.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

omg omg i got the right answer. thank you sooooo much

OpenStudy (phi):

for the quotient rule you should do \[\frac{d}{dx}\left( \frac{x^2}{y^2}\right)= \frac{y^2 \frac{d}{dx}x^2 - x^2 \frac{d}{dx}y^2}{y^4}\]

OpenStudy (phi):

the numerator gives you \[ y^2 \frac{d}{dx}x^2 - x^2 \frac{d}{dx}y^2 \\ y^2 (2x) - x^2 2y y' \\ 2x y^2 - 2x^2 y \frac{x^2}{y^2}\\ 2xy^2 - 2\frac{x^4}{y} \] now divide all that by y^4 \[ \frac{1}{y^4} \left( 2xy^2 - 2\frac{x^4}{y}\right)\\ \frac{2x}{y^2} - \frac{2x^4}{y^5} \]